OpenStudy (anonymous):

Find the x- and y-intercepts of the line y = 4x + 16. x-intercept: (4, 0); y-intercept: (0, 16) x-intercept: (-4, 0); y-intercept: (0, 16) x-intercept: (0, 4); y-intercept: (16, 0) x-intercept: (0, -4); y-intercept: (16, 0)

OpenStudy (anonymous):

so is it the 2nd answer

OpenStudy (anonymous):

Yes. You can tell that from the equation, as when X=0, that is the Y intercept (i.e. y=4(0)+16=16), and when Y=0 that is the X intercept (i.e. 0=4x+16, -4x=16, x=-4)

OpenStudy (anonymous):

how do you do it tho if their isent really an equation for example

OpenStudy (anonymous):

Find the x- and y-intercepts of the line y = 20x. x-intercept: (0, 0); y-intercept: (0, 0) x-intercept: (-20, 0); y-intercept: (0, 20) x-intercept: (-20, 0); y-intercept: (0, 0) x-intercept: (20, 0); y-intercept: (0, -20)

OpenStudy (anonymous):

Same method- make x=0 for the y-intercept, and make y=0 for the x-intercept

OpenStudy (anonymous):

would the answer be the first one?

OpenStudy (anonymous):

Yes